An official from the securities commission estimates that 75% of all investment bankers have profited from the use of insider information. Assume that 15 investment bankers are selected at random from the commission's registry. Using the Binomial Distribution, answer the following questions:
a. Find the probability that at most 10 have profited from insider information.
b. Find the probability that at least 6 have profited from insider information.
c. Find the probability that all 15 have profited from insider information.
d. What is the expected number of investment bankers who have profited from the use of insider information?
e. Find the variance and standard deviation of the number of investment bankers who have profited from the use of insider information.

Answer by stanbon(75887)

An official from the securities commission estimates that 75% of all investment bankers have profited from the use of insider information. Assume that 15 investment bankers are selected at random from the commission's registry. Using the Binomial Distribution, answer the following questions:
--------------------------
Binomial Problems with n = 15 and p=0.75
-------
a. Find the probability that at most 10 have profited from insider information.
P(0<= x <= 10) = binomcdf(15,0.75,10) = 0.3135
-----
b. Find the probability that at least 6 have profited from insider information.
binomedf(15,0.75,6) = 0.0042
----------------------------------------
c. Find the probability that all 15 have profited from insider information.
0.75^15 = 0.0134
-------------------
d. What is the expected number of investment bankers who have profited from the use of insider information?
np = 15*0.75 = 11.25
---------------------
e. Find the variance and standard deviation of the number of investment bankers who have profited from the use of insider information.
variance = npq = 15*0.75*0.252.8125
std = sqrt(variance) = 1.6771
--------------------------------
Cheers,
Stan H.